OpenStudy (natsent06):

Christopher's best time in a race is four seconds faster than his second-best time. His second-best time is 1 1/3 times slower than his best time. Find his best time.

OpenStudy (anonymous):

let best time be x and second best time be 7 according to the first condition x - y = 4 according to the second condition y = 3/4 x now you need to solve the equations.

OpenStudy (anonymous):

*the second best time be y (not 7)

OpenStudy (anonymous):

hmm, isn't y>x

OpenStudy (anonymous):

yes you are right ... second best time is x and best time is y so x-y = 4 x = 3/4 y

OpenStudy (anonymous):

think its x = 4/3 y